Each of the following mortgage loans is for $155,000. The closing costs on each loan will be $4,200, and all points will be paid by the buyer. Which loan requires the most money to be paid up front at closing?

1. Adjustable-rate mortgage: 5 percent interest with 2 points

2. FHA-insured mortgage: 7 percent interest with 1 point

3. Conventional mortgage: 8 percent interest with 0 points

4. All require the same amount at closing.
Answer #1

Answer:

Correct answer is:

1. Adjustable-rate mortgage: 5 percent interest with 2 points

Explanation:

Amount to be paid upfront at closing:

1. Adjustable-rate mortgage = 4200 + 155000 * 2% = $7,300

2. FHA-insured mortgage = 4200 + 155000 * 1% = $5750

3. Conventional mortgage = 4200 + 155000 * 0% = $4,200

Hence option 1 is correct and other options 2, 3 and 4 are incorrect.